# Nagaoka Ferromagnetism in $3\times 3$ Arrays and Beyond

### v1

@@ III.2 $N\times N$ arrays
NF was first predicted for a 2D extended bulk system of electrons with infinite interaction $U$. The bulk limit should be reached as the size of an $N\times N$ array increases. To see if NF for one hole in a half-filled band still occurs for small $t/U$ as N increases and, if so, how the transition point changes as the array size increases, and to see if ferromagnetism can appear for smaller band fillings for larger N we have studied $N\times N$ arrays for N = 4, 5, 6, and 7.
@@
Figure 5: NF for $N\times N$ arrays. (a) NF for $4\times 4$ and $5\times 5$ arrays. Energy of the ground state for each spin $z$ subspace plotted as a function of $|t/U|$, offset by the energy of the maximum-total-spin state. The black dashed line indicates the transition for NF. Both system are in the one hole case, with $S_{max}$ for $4\times 4$ arrays $S_{max}^{4\mathrm{x}4}=15/2$, and for $5\times 5$ arrays $S_{max}^{5\mathrm{x}5}=12$. Here only $s_{z}$ subspaces accessible with ED are shown. (b) Transition point vs N, including both ED and DMRG results. Fitted exponential and power law functions shows how the system approaches the bulk limit (Nagaoka limit $U\rightarrow\infty$) as N gets larger.
Simulation done for $4\times 4$ and $5\times 5$ arrays for one-hole in a half-filled band are shown in Fig. 5a. The larger $N\times N$ arrays behave similarly to the $3\times 3$ and $2\times 2$ arrays, with the transition point decreasing as N gets larger, as shown in Fig. 5b. As the $N\times N$ array gets larger, the onsite interaction $U$ needs to be larger for NF to occur. This agrees with Nagaoka’s original theorem, that ferromagnetism in an infinite lattice occurs for infinite onsite interaction $U$. Fig. 5a includes the results from ED and DMRG (density matrix renormalization group) simulations. For smaller $s_{z}$ the large number of possible spin configurations makes ED difficult even for $4\times 4$ and $5\times 5$ arrays. When ED calculations are not possible, we perform DMRG simulations. Here, we flatten the 2D arrays into 1D chains with long-range hoppings, and obtain their ground state energies through the well-established two-site iterative variational searches [22, 23]. The DMRG calculations confirm the prediction that for larger $N\times N$ arrays, NF exists for a smaller region of $|t/U|$. The dependence of the transition ratio on array size N shown in Fig. 5b out to the $7\times 7$ array has been well fitted to both an exponential and inverse power law on N. In the appendix, a scaling argument is used to support an inverse power law dependence on N. The $7\times 7$ is the first lattice with more internal sites than edge sites. Our results suggest that the bulk limit for NF ferromagnetism with $U$ becoming very large is being approached by N=7. Table 1 summarizes the results for $N\times N$ arrays.
@@
$4\times 4$ [Uncaptioned image] yes 15 $0.0058$ ED, DMRG
$5\times 5$ [Uncaptioned image] yes 24 $0.0023$ ED, DMRG
@@
$7\times 7$ [Uncaptioned image] yes 48 $0.0003$ DMRG
Table 1: $N\times N$ arrays results: geometry, sketch, whether Nagaoka ferromagnetism exists, number of electrons, and transition point, calculational methods.
IV Robustness of NF in $3\times 3$ arrays
In actual experiments with dopant arrays [19], the arrays are not perfectly ordered. Disorder in dopant position and dopant number exists, which can affect the tunnelling and onsite energies. In dopant quantum dot systems intended to have only one dopant at each site, there can be more than one dopant clustered near a site. The properties of that disordered site are determined by the multidopant cluster. This results in a different on-site energy for that particular site compared to the on-site energy that is common to the array sites with a single dopant. It is harder for electrons to hop on and off the disordered site, lowering the participation of that site in the array states. When the energy of the disordered site is sufficiently far off-resonance from the other sites, the disordered site is effectively removed from the array. To investigate the robustness of NF in realistic $3\times 3$ arrays, we consider $3\times 3$ arrays with one site missing to see if ferromagnetism survives or disappears, and, if ferromagnetism survives, what electron fillings display the ferromagnetism. This will provide us insight into how and when NF exists in small finite systems with realistic conditions. For the $3\times 3$ array, there are three different types of site, the corner, edge, and center sites, corresponding to having 2, 3, or 4 hoppings connecting it to other other sites. Removing one site of each type of atom is shown in Fig. 6. We discuss these three cases in the next three sections. We will also discuss ferromagnetism in other finite structures that can be made by removing similar sets of sites in larger $N\times N$ arrays. We summarize the results for removing sites from the $3\times 3$ array and other extended structures in Table 2.
@@ IV Robustness of NF in $3\times 3$ arrays
Geometry
Sketch
NF/IF?
#$\#e$
Transition
Details
$3\times 3$
NF
8
$0.053$
1-hole NF
no corner
NF
7
$0.0104$
1-hole NF
no edge
no
not connected
no center
IF
3
$0.131$
$3e$ IF
no same-side corner
no
not connected
no opposite corner (2 corner-sharing sq)
IF
5
$0.0127$
2-hole IF
3 corner-sharing sq
IF
8
$0.0015$
2-hole IF
4 corner-sharing sq
IF
11
$0.0011$
2-hole IF
5 corner-sharing sq
IF
14
$0.00022$
2-hole IF
2x3 (2 side-sharing sq)
NF
5
$0.019$
1-hole NF
2x4 (3 side-sharing sq)
NF
7
$0.012$
1-hole NF
2x5 (4 side-sharing sq)
NF
9
$0.0082$
1-hole NF
N-site loops
IF
3
$0.017N$ $-0.002$
$3e$ IF
Table 2: Summary table for other geometries, their sketch, whether Nagaoka ferromagnetism (NF) or intinerant ferromagnetism (IF) exists, number of electrons, and transition point if applicable.
IV.1 Removing a Corner
We consider first removing one corner atom (shown in Fig. 6a), which removes the fewest hoppings and gives a structure most similar to the original $3\times 3$ array. After one of the corner atoms is removed, the structure becomes three $2\times 2$ squares stacked together, sharing common sides. NF still exists but now for one-hole in the new half-filled band, i.e. now for 7 electrons. The transition point at $|t/U|\approx 0.0104$ is slightly lower than for the full $3\times 3$ array at $0.0144$. Removing one site can be modeled by increasing the onsite energy of that site. When the onsite energy is zero (same as the other sites in the array), the system is the full $3\times 3$ array. When the onsite energy is infinite, the array has one less site. In this paper, we have discussed these two limits for the onsite energy pf the disordered site. In an upcoming paper, we will show how the ferromagnetism evolves from NF on a $3\times 3$ array to NF on an array with one less corner site and one less electron when the disorder in the site-energy at a corner is increased.
There are two different ways to remove two corner sites in a $3\times 3$ array, as shown in lines 5 and 6 in Table 2. If two corner atoms on the same side are removed, the structure becomes a 2x3 rectangle connected by one extra hopping to an edge site. This geometry does not exhibit GS ferromagnetism for one hole in the half-filled band.
Tasaki showed [2] that for NF to exist the many-body wavefunction must satisfy a connectivity condition. Here, being a connected many-body wavefunction requires that each site be connected to other sites in the lattice by hopping and, more importantly, that any configuration of spins in the array can be connected, via hopping without double occupancy, to any other configuration of the same spins in the array. Connectivity is a necessary but not sufficient condition for saturated ferromagnetism. For an array with a hanging single hopping as shown in line 5 of Table 2, a spin on the isolated site connected to the rest of the array by one hopping can be moved off the isolated site only by exchanging with a hole on its only neighbor. However, the spin can’t be moved further to another site in the array unless a second hole is present. This is shown in Fig. 7a. The many-body wavefunction for this geometry is not connected for one-hole in a half-filled band and can’t realize saturated ferromagnetism for that filling. This prohibits NF for one-hole in a half-filled band. However, the many-body wavefunction becomes connected when there are two or more holes in the half-filled band. However, we didn’t find GS saturated ferromagnetism for any number of holes. This shows that the connectivity condition is necessary but not sufficient.
Figure 7: (a) An example showing when the connectivity condition for NF is violated. To exchange the two spins (in the box), a series of nearest neighbor hoppings is not sufficient, as the spin up and spin down electrons must be on the same site which is strongly suppressed by large repulsion $U$. (b) Example of the connectivity condition being satisfied for 3 electrons on a loop. For any sized loop, 3 electrons is the only filling where the wavefunction is connected. Any pairs of spins (shown in the box) can be exchanged when the other electrons hopped around the loop.
If the two sites are removed from opposite corners of the $3\times 3$ array, the structure becomes two $2\times 2$ plaquettes connected by sharing a common corner site. The ground state for this structure is ferromagnetic only when there are two holes in a half-filled band of the array. The transition is at $|t/U|\approx 0.0127$ and calculations of the pair correlation function show ferromagnetic behaviour below the transition. We refer to this as the itinerant ferromagnetism (IF) case because it is not restricted to the one-hole in a half-filled band of NF. This ferromagnetism for two-holes in a half-filled band can be extended to longer chains of squares. The ground state of the longer chains of corner-sharing squares is ferromagnetic but only when there are two holes in a half-filled band, and not ferromagnetic for the one-hole case (the NF state). The many-body wavefunction is not connected when there is one-hole in a half-filled band, preventing GS saturated ferromagnetism. The site that the two $2\times 2$ plaquettes share acts as a bottleneck. A hole, on one of the $2\times 2$ plaquettes can’t be used to move a spin around another plaquette that has spins at each site. A second hole is needed. Hence ferromagnetism is possible for 2 holes, as we have found. As before, this is a necessary but not sufficient condition. Saturated ferromagnetism does not exist for more than two holes. As the chain gets longer, the transition point also decreases as happens when arrays increase in size.
@@ IV.1 Removing a Corner
IV.2 Removing an Edge Site
@@
In Fig. 6b, the middle site on an edge is removed. We refer to this removed site as an edge site. This leaves the array with two sites at the corners of the array attached to the 2x3 rectangle by single bonds. For this array, there is no ferromagnetism for any number of holes in the half-filled band, i.e. for any electron filling. As seen for an array with same-side corners removed, the many-body wavefunction is not connected for one-hole in a half-filled band of the array with a missing edge site. Thus, NF is not possible. The many-body wavefunction becomes connected for two holes, but there is still no GS ferromagnetism for any number of holes.
@@
If the single hoppings and the isolated sites are removed, the structure turns into a rectangle with internal side-to-side hoppings, and NF appears again for one hole in the half-filled band of the rectangle. The rectangle can be extended to longer chains of $2\times 2$ plaquettes with each pair of adjacent squares sharing a common side. Again, the transition point decreases as the system gets longer. The results for the rectangles can be seen in line $10-12$ in Table 2.
IV.3 Removing the Center Site
@@
When the center site is removed from the $3\times 3$ array, the array becomes an eight-site loop. The many-body wavefunction for this loop is not connected for one-hole in a half-filled band, as shown by Fig. 7b. As a consequence, NF can’t exist. Our calculations confirm this. Interestingly, the 8-site loop does have a ferromagnetic ground state for filling by 3 electrons. This result exists regardless of the size of the loop, and the transition point increases linearly as the number of sites on the loop increases, shown in Fig. 8. In fact, the many-body wavefunction on the loop is connected for a filling of three electrons. The many-body wavefunction is not connected for any higher electron filling. Thus, a ferromagnetic GS should only be possible for a filling of 3 electrons.
The Nagaoka ferromagnetism of $N\times N$ arrays is very different from the GS ferromagnetism of 3-electron N site loops. Nagaoka ferromagnetism in an $N\times N$ array occurs for one-hole in a half-filled band, i.e. for a number of electrons that increases as N increases. The loop ferromagnetism occurs only for 3 electrons, regardless of the number of loop sites N. The transition $t/U$ decreases as an inverse power law of N in $N\times N$ arrays, requiring larger $U$ to achieve ferromagnetism in a larger array. For a loop, the transition point increases linearly with increasing loop size. In the appendix, the inverse power law for $N\times N$ arrays and the linear dependence for loops is explained based on arguments for how the hopping energy and the repulsive energy of interaction scale with system size. The opposite scaling behavior is related to how the hopping energy scales in $N\times N$ arrays with an almost half-filled band and in loops filled with only 3 electrons with system size. The $t/U$ transition ratio for NF in $2\times 2$ arrays is 0.05 and is smaller for larger arrays. The hopping $t$ is much smaller than $U$ at the transition point for all square arrays. The transition point for itinerant ferromagnetism in loops quickly becomes of order 1 with $t$ similar to $U$, as the loop size increases, for all but the smaller loops. The $2\times 2$ array can also be thought as a 4-site loop. In this case, one-hole in a half-filled band is three electrons, so the filling for ferromagnetism is the same whether the array is considered a $2\times 2$ array or a 4-site loop. There is ambiguity in whether ferromagnetism should be consider Nagaoka ferromagnetism or 3-electron loop ferromagnetism. For larger systems this distinction becomes clear.
@@
The specific cases for pentagons and hexagons were discussed in [14], which described a pattern of ferromagnetism appearing in small pentagonal and hexagonal plaquettes at filling factors of roughly 3/10 and 1/4. However, this density dependence appears to be an artifact due to the different geometries considered. In each case, the ferromagnetism occurs for 3 electron filling. This is not a density effect. This filling for ferromagnetism arises because 3-electron filling is the only filling on loops that allow a connected many-body wavefunction.
@@
Figure 8: Dependence of the transition point $|t/U|$ on loop size. As the loop gets larger, the transition point gets larger. Here only loop sizes accessible with ED are shown. Linear fit is shown.

### v2

@@ III.2 $N\times N$ arrays
NF was first predicted for a 2D extended bulk system of electrons with infinite interaction $U$. The bulk limit should be reached as the size of an $N\times N$ array increases. To see if NF for one hole in a half-filled band still occurs for small $t/U$ as $N$ increases and, if so, how the transition point changes as the array size increases, and to see if ferromagnetism can appear for smaller band fillings for larger $N$ we have studied $N\times N$ arrays for $N$ = 4, 5, 6, 7 and 8.
@@
Figure 5: NF for $N\times N$ arrays. (a) NF for $4\times 4$ and $5\times 5$ arrays. Energy of GS for each $s_{z}$ subspace plotted as a function of $|t/U|$, offset by the energy of the maximum $s_{z}$ state. The black dashed line indicates the transition for NF. Both system are in the one hole case, $4\times 4$ arrays have $S_{max}^{4\mathrm{x}4}=15/2$, and for $5\times 5$ arrays $S_{max}^{5\mathrm{x}5}=12$. Here $s_{z}$ subspaces accessible with ED are shown as solid lines, DMRG as dashed lines. (b) Transition point v.s. $N$, including both ED and DMRG results. Fitted power law functions shows how the system approaches the bulk limit (Nagaoka limit $U\rightarrow\infty$) as $N$ gets larger.
Simulation done for $4\times 4$ and $5\times 5$ arrays for one-hole in a half-filled band are shown in Fig.5a. The larger $N\times N$ arrays behave similarly to the $3\times 3$ and $2\times 2$ arrays, with the transition point decreasing as N gets larger, as shown in Fig. 5b. As the $N\times N$ array gets larger, the on-site interaction $U$ needs to be larger for NF to occur. This agrees with Nagaoka’s original theorem, that ferromagnetism in an infinite lattice occurs for infinite on-site interaction $U$. Fig. 5a includes the results from ED and density matrix renormalization group (DMRG) simulations. For smaller $s_{z}$ the large number of possible spin configurations makes ED difficult even for $4\times 4$ and $5\times 5$ arrays. When ED calculations are not possible, we perform DMRG simulations. Here, we flatten the 2D arrays into 1D chains with long-range hoppings, and obtain their ground state energies through the well-established two-site iterative variational searches [27, 28]. The DMRG calculations confirm the prediction that for larger $N\times N$ arrays, NF exists for a smaller region of $|t/U|$. The dependence of the transition ratio on array size $N$ shown in Fig. 5b out to the $8\times 8$ array has been well fitted to an inverse power law on $N$. In the Appendix A, a scaling argument is used to support an inverse power law dependence on $N$. The $7\times 7$ is the first lattice with more internal sites than edge sites. Our results suggest that the bulk limit for NF ferromagnetism with $U$ becoming very large is being approached by $N=8$. Table 1 summarizes the results for $N\times N$ arrays.
@@
$4\times 4$ [Uncaptioned image] yes 15 $0.0061$ ED, DMRG
$5\times 5$ [Uncaptioned image] yes 24 $0.0024$ ED, DMRG
@@
$7\times 7$ [Uncaptioned image] yes 48 $0.00055$ DMRG
$8\times 8$ [Uncaptioned image] yes 63 $0.00023$ DMRG
Table 1: $N\times N$ arrays results: geometry, sketch, whether Nagaoka ferromagnetism exists, number of electrons, transition point, and calculational methods.
IV Different Geometries beyond the $3\times 3$ Arrays
The sites in a semiconductor array need not be identical. The site energies can be controlled by bias. In dopant systems, disorder in dopant position and dopant number can affect the tunnelling and on-site energies. For example, there can be more than one dopant clustered near a site. The properties of that disordered site will be determined by the multidopant cluster which will have a different on-site energy. It is harder for electrons to hop on and off a disordered site with a shifted on-site energy, lowering the participation of that site in the array states. When the energy of the disordered site is sufficiently far off-resonance from the other sites, the disordered site is effectively removed from the array. To investigate the robustness of NF in fabricated $3\times 3$ arrays with disordered sites, we consider $3\times 3$ arrays with one site missing to see if ferromagnetism survives or disappears, and, if ferromagnetism survives, what electron fillings display the ferromagnetism. This will provide us insight into how and when NF exists in small finite systems with variations in on-site energies. In this paper we look at the limit when the on-site energy difference is large enough to remove the site. In an upcoming paper, we will show how the ferromagnetism evolves as on-site energy differences vary.
@@ IV Different Geometries beyond the $3\times 3$ Arrays
For the $3\times 3$ array, there are three different types of site, the corner, edge, and center sites, corresponding to having 2, 3, or 4 hoppings connecting it to other other sites. Removing one site of each type of atom is shown in Fig. 6. We discuss these three cases in the next three sections. We also discuss ferromagnetism in other finite structures that can be made by removing similar sets of sites in larger $N\times N$ arrays. We summarize the results for removing sites from the $3\times 3$ array and other extended structures in Table 2.
Geometry Sketch NF/IF? #$\#e$ Transition Details
$3\times 3$ [Uncaptioned image] NF 8 $0.053$ 1-hole NF
no corner [Uncaptioned image] NF 7 $0.0104$ 1-hole NF
no edge [Uncaptioned image] no not connected
no center [Uncaptioned image] IF 3 $0.131$ $3e$ IF
no same-side corner [Uncaptioned image] no not connected
no opposite corner (2 corner-sharing sq) [Uncaptioned image] IF 5 $0.0127$ 2-hole IF
2x3 (2 side-sharing sq) [Uncaptioned image] NF 5 $0.019$ 1-hole NF
$N$-site loops [Uncaptioned image] IF 3 $0.017N$ $-0.002$ $3e$ IF
Table 2: Summary table for other geometries, their sketch, whether Nagaoka ferromagnetism (NF) or itinerant ferromagnetism (IF) exists, number of electrons, and transition point if applicable.
IV.1 Removing Corner Atoms - Rectangular Arrays
IV.1.1 Removing One Corner Atom
We consider first removing one corner atom (shown in Fig. 6a), which removes the fewest hoppings and gives a structure most similar to the original $3\times 3$ array. After one of the corner atoms is removed, the structure becomes three $2\times 2$ squares stacked together, sharing common sides. NF still exists but now for one-hole in the new half-filled band, i.e. now for 7 electrons. The transition point at $|t/U|\approx 0.0104$ is slightly lower than for the full $3\times 3$ array at $0.0144$. Removing one site can be modeled by increasing the on-site energy of that site. When the on-site energy is zero (same as the other sites in the array), the system is the full $3\times 3$ array. When the on-site energy is infinite, the array has one less site. In this paper, we discuss these two limits for the on-site energy of the disordered site. In an upcoming paper, we will show how the ferromagnetism evolves from NF on a $3\times 3$ array to NF on an array with one less corner site and one less electron as the energy shift of the disordered corner site is increased.
IV.1.2 Removing Two Corners on the Same Side
There are two different ways to remove two corner sites in a $3\times 3$ array, as shown in lines 5 and 6 in Table 2. If two corner atoms on the same side are removed, the structure becomes a $2\times 3$ rectangle connected by one extra hopping to an edge site. This geometry does not exhibit GS ferromagnetism for one hole in the half-filled band.
IV.1.3 Role of Wavefunction Connectivity
Figure 7: (a) An example showing when the connectivity condition for NF is violated. To exchange the two spins (in the box), a series of nearest neighbor hoppings is not sufficient, as the spin up and spin down electrons must be on the same site which is strongly suppressed by large repulsion $U$. (b) Example of the connectivity condition being satisfied for 3 electrons on a loop. For any sized loop, 3 electrons is the only filling where the wavefunction is connected. Any pair of spins (shown in the box) can be exchanged when all electrons are hopped around the loop.
Tasaki showed [2] that for NF to exist the many-body wavefunction must satisfy a connectivity condition. Here, being a connected many-body wavefunction requires that each site be connected to other sites in the lattice by tunnel coupling and, more importantly, that any configuration of spins in the array can be connected, via hopping without double occupancy, to any other configuration of the same spins in the array. Connectivity is a necessary but not sufficient condition for saturated ferromagnetism. For an array with a hanging single hopping as shown in line 5 of Table 2, a spin on the isolated site connected to the rest of the array by one hopping can be moved off the isolated site only by exchanging with a hole on its only neighbor. However, the spin can’t be moved further to another site in the array unless a second hole is present. This is shown in Fig. 7a. The many-body wavefunction for this geometry is not connected for one-hole in a half-filled band and can’t realize saturated ferromagnetism for that filling. This prohibits NF for one-hole in a half-filled band. However, the many-body wavefunction becomes connected when there are two or more holes in the half-filled band. However, we don’t find GS saturated ferromagnetism for any number of holes. This shows that the connectivity condition is necessary but not sufficient.
IV.1.4 Removing Two Opposite Corners
If the two sites are removed from opposite corners of the $3\times 3$ array, the structure becomes two $2\times 2$ plaquettes connected by sharing a common corner site. The ground state for this structure is ferromagnetic only when there are two holes in a half-filled band of the array, that is, when there are five electrons. The transition is at $|t/U|\approx 0.0127$ and calculations of the pair correlation function show ferromagnetic behaviour below the transition. We refer to this as the itinerant ferromagnetism (IF) case because it is not possible for one-hole doping of a half-filled band as in NF. This ferromagnetism for two-hole doping in a half-filled band can be extended to longer chains of squares. The ground state of the longer chains of corner-sharing squares is ferromagnetic but only when there are two holes in a half-filled band, and not ferromagnetic for the one-hole case (the NF state). The many-body wavefunction is not connected when there is one-hole in a half-filled band, preventing GS saturated ferromagnetism. The site shared by the two $2\times 2$ plaquettes acts as a bottleneck. A hole on one of the $2\times 2$ plaquettes can’t be used to move a spin around another plaquette that has spins at each site. A second hole is needed. Hence ferromagnetism is possible for two holes, as we have found. As before, this is a necessary but not sufficient condition. Saturated ferromagnetism does not exist for more than two holes. As the chain gets longer, the transition point also decreases as happens when arrays increase in size as seen in Fig 8. (For details, see Appendix B.2).
Figure 8: Two-hole ferromagnetism in a chain with $N$ $2\times 2$ plaquettes sharing corners on a common diagonal. Transition point versus the number $N$ of plaquettes. The fitted power law function is shown as a dotted line.
@@ IV.1.4 Removing Two Opposite Corners
IV.2 Removing an Edge Site - One-fold Coordinated Sites
@@
When the middle site on an edge is removed, here referred to as an edge site, the array has two corner sites attached to a $2\times 3$ rectangle by single bonds. The corner sites are now one-fold coordinated. For this array, there is no ferromagnetism for any number of holes in the half-filled band, i.e. for any electron filling. As seen for an array with same-side corners removed, the many-body wavefunction is not connected for one-hole doping of the half-filled band of the array with a missing edge site. Thus, NF is not possible. The many-body wavefunction becomes connected for two holes, but there is still no GS ferromagnetism for any number of holes.
@@
If the single hoppings and the isolated sites are removed, the structure turns into a rectangle with internal side-to-side hoppings, and NF appears again for one hole in the half-filled band of the rectangle. The rectangle can be extended to longer chains of $2\times 2$ plaquettes with each pair of adjacent squares sharing a common side. Again, the transition point decreases as the system gets longer. The results for the rectangles are discussed in Appendix B.1.
IV.3 Removing the Center Site - Loops
IV.3.1 Three-electron Loop Ferromagnetism
@@
When the center site is removed from the $3\times 3$ array, the array becomes an eight-site loop. The many-body wavefunction for this loop is not connected for one-hole in a half-filled band, as shown by Fig. 7b. As a consequence, NF can’t exist. Our calculations confirm this. Interestingly, the 8-site loop does have a ferromagnetic ground state for filling by 3 electrons. This result exists regardless of the size of the loop, and the transition point increases linearly as the number of sites on the loop increases, shown in Fig. 9. In fact, the many-body wavefunction on the loop is connected for a filling of three electrons. The many-body wavefunction is not connected for any higher electron filling. Thus, a ferromagnetic GS should only be possible for a filling of 3 electrons or less.
@@
The specific cases for pentagons (five-site loops) and hexagons (six-site loops) were discussed in [16], which described ferromagnetism appearing in small pentagonal and hexagonal plaquettes at filling factors of roughly 3/10 and 1/4. However, this density dependence appears to be an artifact due to the different geometries considered. In each case, the ferromagnetism occurs for 3 electron filling. This is not a density effect. This filling for ferromagnetism arises because three-electron filling is the only filling on loops that allows a connected many-body wavefunction.
The Nagaoka ferromagnetism of $N\times N$ arrays is very different from the GS ferromagnetism of three-electron $N$-site loops. Nagaoka ferromagnetism in an $N\times N$ array occurs for one-hole in a half-filled band. The number of electrons in the NF state increases as $N$ increases. However, for any $N$, the NF state has one hole. In contrast, loop ferromagnetism occurs only for 3 electrons, regardless of the number of loop sites $N$. The transition $t/U$ decreases as an inverse power law of $N$ in $N\times N$ arrays, requiring larger $U$ to achieve ferromagnetism in a larger array. For a loop, the transition point increases linearly with increasing loop size.
@@
Figure 9: Dependence of the transition point $|t/U|$ on loop size. As the loop gets larger, the transition point gets larger. Here only loop sizes accessible with ED are shown. A linear fit is shown.
These very different scalings are determined in part by the very different band fillings: 3 electron filling for loops and one-hole in a half-filled band for NF. The loop ferromagnetism occurs for a system with a few (i.e. three) electrons in an empty band. NF occurs for a few holes (i.e. one) in a half-full band. Explaining the different scalings for loop ferromagnetism and NF starts with this difference in band filling.
IV.3.2 Scaling Arguments for the Transition to Ferromagnetism
In the following, we will lay out the basic arguments for the two scaling relations to show why they are so different. Additional details and comments are provided in Appendix A. The linear dependence for loops is explained based on how the hopping energy and the repulsive energy of interaction scale with loop length. For three-electron filling, only states near the bottom of the band are occupied unless $N$ is small. Due to the confinement to the loop, single-particle states near the bottom of the band have energy shifts above the band edge that vary with loop size as $1/N^{2}$. As a consequence, the kinetic energy difference between the ferromagnetic state and the non-ferromagnetic state should vary with increasing loop length as $t/N^{2}$. The kinetic energy of the ferromagnetic state, with all three electrons with the same spin will be higher than the kinetic energy of the non-ferromagnetic state with both majority and minority spin electrons.
The crossover from ferromagnetic to not ferromagnetic occurs when the kinetic energy cost to be ferromagnetic is compensated by the reduction in the repulsive interaction energy when the spins are aligned. There is no interaction energy in the ferromagnetic state because all electrons have the same spin. In the non-ferromagnetic state, there is an interaction energy $U$ for each pair of opposite spin electrons on the same site. The probability that two electrons of opposite spin are on the same site in the loop varies with loop length as $1/N$ and, as a result, the interaction energy varies as $U/N$. The $t/U$ for the transition between the ferromagnetic and non-ferromagnetic states occurs when the kinetic energy cost that varies as $t/N^{2}$ is compensated by the repulsive interaction energy that varies as $U/N$. The $t/U$ at the transition should scale as $N$ as the loop size varies. This linear scaling for loop ferromagnetism is determined by the three electron filling and the scaling of the kinetic and interaction energies.
The arguments to support the $1/N^{3.5}$ size-scaling of the transition $t/U$ for NF in $N\times N$ arrays with one hole in half-filled band are more complicated. Here we summarize the key results. The details are given in Appendix A. The arguments are easiest to develop in the infinite $U/t$ limit. The transition is identified by why and when the excited, non-ferromagnetic states cross the ferromagnetic ground states as $U$ decreases. In the infinite $U$ limit, the energy levels of states with one hole in a half-filled band can be determined by lowest order degenerate perturbation theory where $|t/U|$ is the small perturbation parameter. In the infinite $U$ limit, the fully polarized, single-hole state is the ground state with the energy for one hole hopping freely in the array. Low-lying partially polarized excited states are shifted up from the ground state by energies that scale between $|t|/N^{3}$ and $|t|/N^{4}$ due to how the hole is localized by scattering from minority spins in the array. The fully polarized ground state in the infinite $U$ limit has no repulsive interaction energy so its energy remains fixed as $U$ becomes finite. However, the higher lying, partially polarized states gain exchange energy $-t^{2}/U$ that lowers their energy, because hopping to double occupied sites becomes possible for finite $U$. The transition to a non-ferromagnetic state occurs when the energy gain due to exchange lowers the excited state energies enough to cross the fully polarized level. This occurs at a $|t|/U$ that scales between $|t|/N^{3}$ and $|t|/N^{4}$. This is consistent with the $1/N^{3.5}$ scaling obtained from the computational results.
IV.3.3 Stark Contrast between Loop and Nagaoka Ferromagnetism
The scaling arguments point to the very different nature of loop and Nagaoka ferromagnetism. Linear scaling of the transition $|t/U|$ for loop ferromagnetism occurs when the kinetic energy cost for being ferromagnetic is compensated by the reduced repulsive interaction when ferromagnetic. For Nagaoka ferromagnetism in $N\times N$ arrays, the ferromagnetic phase has lower hopping energy making it the ground state in the infinite $U/t$ limit. The not-ferromagnetic phases are lowered in energy due to exchange interactions that couple the not-ferromagnetic states to states that have doubly-occupied sites. This contrasting behavior is connected to loop ferromagnetism occurring at low electron filling and Nagaoka ferromagnetism occurring at single-hole doping.
The $t/U$ transition ratio for NF in $2\times 2$ arrays is $0.053$ and is smaller for larger arrays. The hopping $t$ is much smaller than $U$ at the transition point for all square arrays. The transition point for itinerant ferromagnetism in loops quickly becomes of order 1 with $t$ similar to $U$, as the loop size increases, for all but the smallest loops. This provides another stark contrast between loop and Nagaoka ferromagnetism.
There is one case where loop and Nagaoka ferromagnetism may be the same. The $2\times 2$ array can also be thought as a 4-site loop. In this case, one hole in a half-filled band is three electron filling, so the filling for ferromagnetism is the same whether the array is considered a $2\times 2$ array or a 4-site loop. There is ambiguity in whether ferromagnetism should be consider Nagaoka ferromagnetism or three-electron loop ferromagnetism. For larger systems this distinction becomes clear.
IV.3.4 Particle Correlation in Loop Ferromagnetism